OpenStudy (anonymous):

\[-1.156+n/3.4=-.126\]

OpenStudy (anonymous):

adding both sides by 1.156 \[-1.156+\frac{n}{3.4}+1.156=-.126+1.156\] \[\frac{n}{3.4}=1.03\] multiply both sides by 3.4 \[\frac{n}{3.4}*3.4=1.03*3.4\] \[n=3.502\]